### v2

# All loop soft photon theorems and higher spin currents on the celestial sphere

Abstract
Soft factorization theorems can be reinterpreted as Ward identities for (asymptotic) symmetries of scattering amplitudes in asymptotically flat space-time. In this paper we study the symmetries implied by the all loop soft photon theorems when all external particles are massless. Loop level soft theorems are qualitatively different from the tree level soft theorems because loop level soft factors contain multi-particle sums. If we want to interpret them as Ward identities then we need to introduce additional fields which live on the celestial sphere but do not appear as asymptotic states in any scattering experiment. For example, if we want to interpret the one-loop exact $\mathcal{O}(\ln\omega)$ soft theorem for a positive helicity soft photon (with energy $\omega$) as a Ward identity then we need to introduce a pair of antiholomorphic currents on the celestial sphere which transform as a doublet under the $SL(2,\mathbb{R})_{R}$. We call them dipole currents because the corresponding charges measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere. More generally, the soft photon theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ for every $j\in\frac{1}{2}\mathbb{Z}_{+}$ gives rise to $(2j+1)$ antiholomorphic currents which transform in the spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. These currents exist in the quantum theory because they follow from loop level soft theorems. We argue that under certain circumstances the (classical) algebra of the higher spin currents is the wedge subalgebra of the $w_{1+\infty}$.
I Introduction
Soft factorization theorems can be reinterpreted as Ward identities for symmetries of scattering amplitudes in asymptotically flat space-time. This plays an important role in Celestial Holography Strominger:2017zoo ; Pasterski:2016qvg ; Banerjee:2018gce which posits that the holographic dual of quantum theory of gravity in asymptotically flat four dimensional space time is a CFT2 living on the celestial two-sphere. The symmetries of the scattering amplitudes arising out of soft theorems are global symmetries of the Celestial CFT2 and in many cases the global symmetry is infinite dimensional Strominger:2013lka ; Barnich:2009se ; Banerjee:2020zlg ; Guevara:2021abz ; Strominger:2021mtt ; Himwich:2021dau ; Adamo:2021lrv ; Freidel:2021ytz ; Stieberger:2023fju ; Melton:2024akx . This is a powerful statement because one can constrain four dimensional scattering amplitudes Banerjee:2020zlg ; Banerjee:2023zip ; Fan:2022vbz ; Costello:2023vyy using the infinite dimensional global symmetries. Moreover, understanding how these symmetries are realized in celestial CFT2 could shed light on the connection between celestial holography and string theory in asymptotically flat spacetime Kervyn:2025wsb ; Donnay:2023kvm ; Banerjee:2025tec .
In this paper, we study the symmetry interpretation of the soft photon theorems that remain valid even at loop level. Consequently, the symmetries implied by these soft theorems are genuine symmetries of the holographic dual celestial CFT2, persisting even after quantum corrections are taken into account. It is well known that the leading soft photon theorem, derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx , is universal and exact at tree level, receiving no loop corrections (up to a subtlety in pQCD with massless quarks as pointed out in Ma:2023gir ). However, since four-dimensional $S$-matrices are infrared divergent due to the long-range interactions mediated by massless particles, deriving finite subleading corrections has long remained a challenge Bern:2014oka ; He:2014bga . By interpreting soft factors as ratios of $S$-matrices with and without an external soft photon, and by systematizing the exponentiation of infrared divergences, it is possible to extract well-defined subleading soft theorems from the IR-finite parts of the ratios of $S$-matrices. This program was successfully carried out in Sahoo:2018lxl , with a proof of universality subsequently established in Krishna:2023fxg . The resulting subleading soft photon theorem appears at $\mathcal{O}(\ln\omega)$ in the soft photon energy $\omega$ and is one loop exact. The $\mathcal{O}(\ln\omega)$ soft factor consists of two distinct contributions, termed the “classical” and “quantum” parts in Sahoo:2018lxl . The classical contribution encodes information about the classical electromagnetic waveform in the low-frequency expansion, as demonstrated through independent classical analyses in Laddha:2018myi ; Saha:2019tub . By contrast, a clear physical interpretation of the quantum contribution remains elusive to date. An intriguing feature of the subleading soft photon theorem is that the classical contribution vanishes for massless charged particles, while the quantum contribution remains non-zero. This observation motivates a deeper investigation of the quantum part of subleading soft photon theorem in massless QED. In particular, we aim to understand it as a Ward identity in the framework of celestial CFT and to uncover the associated symmetry structure.
From a four-dimensional bulk perspective, the classical part of the subleading soft photon theorem at $\mathcal{O}(\ln\omega)$ has already been understood as a Ward identity associated with a large gauge symmetry. [^1] The Celestial OPE of gluons and gravitons at one-loop were studied in Bhardwaj:2022anh ; Krishna:2023ukw . This symmetry acts on charged matter fields as a local enhancement of global phase rotations, referred to as “superphaserotations” in Campiglia:2019wxe ; AtulBhatkar:2019vcb ; Choi:2025mzg . In the present work we provide an interpretation of the quantum contribution to the $\mathcal{O}(\ln\omega)$ soft photon theorem in massless QED as a symmetry of the dual Celestial CFT2. Furthermore, we analyze higher subleading corrections to the soft photon theorem at orders $\omega^{n}(\ln\omega)^{n+1}$ as conjectured and explicitly derived up to sub-subleading order in Sahoo:2020ryf ; Karan:2025ndk , and argued to be $(n+1)$-loop exact. Analogously to the subleading soft photon theorem, we argue that at higher subleading orders only the quantum contributions survive in massless QED.
Loop level soft theorems are qualitatively different from the tree level ones. It turns out that if we want to interpret them as Ward identities of symmetries in a Celestial CFT2 then we need to introduce additional fields on the celestial sphere which do not appear as asymptotic states in scattering experiments. In particular, they are not related to the soft particle in any obvious manner. Now since there are an infinite number of loop level soft photon theorems we get an infinite number of chiral currents in this way. For every $j\in\frac{1}{2}\mathbb{Z_{+}}$ the soft theorem at $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ gives rise to $(2j+1)$ chiral currents which transform in the finite dimensional spin-$j$ representation of the $SL(2,\mathbb{R})_{R}$. The charges corresponding to the $j=1/2$ currents measure the monopole and the dipole moment of an electrically charged particle on the celestial sphere and hence called a dipole current. The other spin-$j$ currents can be constructed by taking the (normal ordered) product of $2j$ dipole currents. Depending on the value of a real parameter $k$ which appears in the OPE of the dipole currents, the algebra of the higher spin currents can be Abelian ($k=0$) or the wedge subalgebra ($k<<1$) of the $w_{1+\infty}$.
Appearance of higher spin currents is not new in the context of Celestial holography. Both soft gluons and gravitons generate infinite dimensional higher spin algebras known as the $S$-algebra and the wedge subalgebra of $w_{1+\infty}$ Strominger:2021mtt ; Adamo:2021lrv ; Freidel:2021ytz ; Himwich:2021dau . However, the infinite number of higher spin currents we get from the loop level soft photon theorems and the symmetries they generate are robust against quantum corrections. It is to been how one can exploit this feature to constrain scattering amplitudes or relate the Celestial holography to string theory in asymptotically flat space-time [^2] We have more to say on this in an upcoming paper..
The paper is organized as follows. In sections II and III we discuss the loop level soft theorems and their representation in the conformal basis. In section-IV we introduce the dipole current in order to write down the one-loop exact soft photon theorem as a Ward-identity. In the following section till section-VII we discuss conformal transformation property of the dipole operator and correlation functions with multiple insertions of dipole operators. In section- VIII we discuss soft photon theorems beyond one loop and show how they give rise to higher spin currents. In section-IX we discuss the algebra of the higher spin currents and point out the role of $k$ in determining the algebra. In section-X we discuss the issue of the closure of the algebra for finite value of $k$ and the possibility of the algebra getting enhanced to the wedge subalgebra of $W_{1+\infty}$. In section-XI we discuss potential problems which arise in determining the value of $k$.
II All loop soft theorems in momentum space
Consider the scattering of massless charged particles in a generic theory of quantum electrodynamics in four spacetime dimensions, where, apart from electromagnetism, no other long-range interaction is present. The soft photon theorem valid at all order in perturbation theory relating the scattering amplitude $\mathcal{A}^{(\text{em})}_{N+1}$, involving $N$ number of finite-energy charged particles with charges $\{e_{a}\}$ and momenta $\{p_{a}\}$, and one outgoing soft photon with polarization $\varepsilon^{h}_{\mu}$, helicity $h$, and momentum $k^{\mu}=-\omega\mathbf{n}^{\mu}$ (where $\omega$ is the photon energy and $\mathbf{n}^{\mu}$ is the unit null vector specifying the emission direction), to the scattering amplitude $\mathcal{A}^{(\text{em})}_{N}$ without the external photon reads: [^3] As is well known, in four spacetime dimensions both $\mathcal{A}^{(\text{em})}_{N+1}$ and $\mathcal{A}^{(\text{em})}_{N}$ are infrared (IR) divergent. However, since the soft factor appears as a ratio of these amplitudes, it is IR finite in the sense that the soft theorem in (1) relates the IR-finite hard parts of the two amplitudes, following the arguments of Sahoo:2018lxl ; Arkani-Hamed:2020gyp ; Krishna:2023fxg .
$\displaystyle\mathcal{A}^{\text{(em)}}_{N+1}\overset{\omega\rightarrow 0}{=}S_{\text{em}}\left(\{e_{a},p_{a}\},(\varepsilon^{h},k)\right)\,\mathcal{A}^{\text{(em)}}_{N}\,+\,\mathcal{R}^{\text{(em)}}_{N+1}$, (1)
where $S_{\text{em}}$ represents the universal soft photon factor determined only in terms of the scattering data specified in its argument and $\mathcal{R}^{\text{(em)}}_{N+1}$ represents the non-factorized part of $\mathcal{A}^{\text{(em)}}_{N+1}$ in the soft limit. [^4] For the all-loop resumed S-matrix in a generic theory of QED, $\mathcal{R}^{\text{(em)}}_{N+1}$ cannot be related to $\mathcal{A}^{\text{(em)}}_{N}$, and hence it does not factorize. However, if we restrict to the tree-level S-matrix, then $\mathbb{R}_{\text{em},N+1}^{(0,0)}$ factorizes and yields the well-known Low’s soft photon theorem, with a specific contribution from the non-minimal coupling through the field strength Low:1954kd ; Elvang:2016qvq . At one-loop order, $\mathbb{R}_{\text{em},N+1}^{(0,0)}$ acquires non-universal contributions originating from different regions of the loop diagrams which fails to factorize Krishna:2023fxg . They have the following expansion in the soft limit:
$\begin{gathered}S_{\text{em}}\left(\{e_{a},p_{a}\},(\varepsilon^{h},k)\right)\overset{\omega\rightarrow 0}{=}\sum_{n=-1}^{\infty}\mathbb{S}_{\text{em}}^{(n)}\times\omega^{n}(\ln\omega)^{n+1},\\ \mathcal{R}^{(\text{em})}_{N+1}\overset{\omega\rightarrow 0}{=}\sum_{n=0}^{\infty}\sum_{t=0}^{n}\mathbb{R}_{\text{em},N+1}^{(n,t)}\,\omega^{n}(\ln\omega)^{t}.\end{gathered}$ (2)
The soft factor coefficient $\mathbb{S}_{\text{em}}^{(n)}$ has been argued to be $(n+1)$-loop exact, and is derived in Gell-Mann:1954wra ; Low:1954kd ; Weinberg:1965nx for n=1n=-1, Sahoo:2018lxl for $n=0$ and Sahoo:2020ryf for $n=1$ for scattering of massive charged particles and proved to be universal in Krishna:2023fxg . Below, we present their expressions for the scattering of massless charged particles, together with a conjectured form valid for generic $n$, motivated by the analysis in appendix A and inspired by refs. Sahoo:2020ryf ; Karan:2025ndk .
$\displaystyle\begin{gathered}\mathbb{S}^{(-1)}_{\text{em}}=-\sum_{a=1}^{N}e_{a}\frac{\varepsilon^{h}_{\mu}p_{a}^{\mu}}{p_{a}.\mathbf{n}},\\ \mathbb{S}^{(n)}_{\text{em}}=-\frac{1}{(n+1)!(4\pi^{2})^{n+1}}\sum_{\begin{subarray}{c}a,b=1\\ b\neq a\end{subarray}}^{N}\frac{e^{2}_{a}e_{b}}{p_{a}.p_{b}}\frac{\varepsilon^{h}_{\mu}\mathbf{n}_{\nu}}{p_{a}.\mathbf{n}}\left(p_{a}^{\mu}p_{b}^{\nu}-p_{a}^{\nu}p_{b}^{\mu}\right)\left(\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}\frac{e_{a}e_{c}}{p_{a}.p_{c}}(p_{c}.\mathbf{n})\right)^{n},\end{gathered}$ (5)
for $n\geq 0$. The soft factors $\mathbb{S}^{(n)}_{\text{em}}$ for $n\geq 0$ presented above are intrinsically quantum contributions to the soft photon theorems, since the classical soft photon factors vanish in the massless limit. [^5] The classical contribution to $\mathbb{S}^{(n)}_{\text{em}}$ for the scattering of massive charged particles was derived in Karan:2025ndk . It vanishes in the massless limit, which is physically expected since, even in the presence of a long-range electromagnetic force, massless charged particles always move at the speed of light in flat spacetime, do not accelerate, and therefore do not radiate. The terminology “classical” and “quantum” follows from Sahoo:2018lxl . In the above expressions, we adopt the mostly positive convention for the Minkowski metric. The charges and energies of incoming particles are taken with their physical signs, while those of outgoing particles acquire an additional negative sign, ensuring that charge and momentum conservation are expressed as $\sum\limits_{a}e_{a}=0$ and $\sum\limits_{a}p_{a}=0$, respectively. From the soft expansion (1) and (2), the soft factor coefficient $\mathbb{S}_{\text{em}}^{(n)}$ can be extracted by using the following projection operator
$\mathbb{S}_{\text{em}}^{(n)}=\frac{1}{(n+1)!}\lim_{\omega\rightarrow 0}\frac{d^{n+1}}{d\omega^{n+1}}\omega\prod_{\ell=0}^{n+1}\frac{1}{\ell!}\left(\omega\frac{d}{d\omega}-(\ell-1)\right)^{\ell}\left[\frac{\mathcal{A}^{\text{(em)}}_{N+1}}{{\mathcal{A}^{\text{(em)}}_{N}}}\right$], (6)
where the projection operator for $n=0$ was already found in Campiglia:2019wxe .
In the presence of additional long-range interactions mediated by other massless fields, the soft factor coefficients $\mathbb{S}_{\text{em}}^{(n)}$ receive further universal corrections for all $n\geq 0$. For instance, the correction to $\mathbb{S}_{\text{em}}^{(0)}$ due to gravitational interactions was analyzed in Sahoo:2018lxl . In the context of the results presented above, even if we restrict attention to theories in which the only massless fields are the photon and the charged fields appearing as external hard particles, loop diagrams involving the same massless charged particles, specifically ladder and crossed-ladder topologies, can still generate additional contributions to $\mathbb{S}_{\text{em}}^{(n)}$ when three-point interaction vertices among the massless charged particles are allowed. This effect is discussed in a footnote of Appendix A for the case of scalar charged particle scattering. Consequently, the universality of the soft factors $\mathbb{S}_{\text{em}}^{(n)}$ for $n\geq 0$ is more restrictive than that of the leading soft factor $\mathbb{S}_{\text{em}}^{(-1)}$, as discussed in Krishna:2023fxg .
III Soft theorems in conformal primary basis
In order to interpret the soft theorems (5) as Ward identities in a CFT2 living on the celestial sphere we transform to the conformal primary basis. We parametrize the soft momentum $k^{\mu}$, positive helicity photon polarization vector $\varepsilon^{+\mu}(k)$ and the momentum $p^{\mu}$ of a massless hard particle as
$\displaystyle\begin{gathered}k^{\mu}=\eta\omega\,\mathbf{n}^{\mu}(w,\bar{w}),\ \omega\rightarrow 0+\\ \mathbf{n}(w,\bar{w})=\big(1+w\bar{w},w+\bar{w},-i(w-\bar{w}),1-w\bar{w}\big)\\ \quad\varepsilon^{+\mu}(k)=\frac{1}{\sqrt{2}}(\bar{w},1,-i,-\bar{w})\\ p^{\mu}=\eta\omega\mathbf{n}(z,\bar{z})\end{gathered}$ (11)
where $\eta=\pm 1$ for an incoming (outgoing) particle and $(w,\bar{w})((z,\bar{z}))$ represent the (stereographic) coordinates of the Celestial sphere. [^6] The charges $\{e_{a}\}$ are related to their actual charge values $\{q_{a}\}$ by relation $e_{a}=\eta_{a}q_{a}$. We also define celestial primary field $\phi^{(\eta)}_{h,\bar{h}}(z,\bar{z},e)$ by Mellin transforming the creation and annihilation operators $b^{\dagger}$ and $b$ as Pasterski:2016qvg
$\displaystyle\begin{gathered}\phi^{(+1)}_{h,\bar{h}}(z,\bar{z},e)=\int_{0}^{\infty}d\omega\,\omega^{\Delta-1}\,b^{\dagger}(p,\sigma,e),\\ \phi^{(-1)}_{h,\bar{h}}(z,\bar{z},e)=\int_{0}^{\infty}d\omega\,\omega^{\Delta-1}\,b(p,\sigma,e),\end{gathered}$ (14)
where $\Delta$ is in general a complex number, $\sigma\in\frac{1}{2}\mathbb{Z}$ is the helicity of the particle and $e$ is the electric charge. One can show that under Lorentz transformation which acts on the Celestial sphere as the group ($SL(2,\mathbb{C})$) of global conformal transformations, the field $\phi^{(\eta)}_{h,\bar{h}}(z,\bar{z},e)$ transforms as a conformal primary with holomorphic and anti-holomorphic conformal dimensions given by $h=\frac{1}{2}(\Delta+\sigma)$ and $\bar{h}=\frac{1}{2}(\Delta-\sigma)$, respectively.
Now after transforming to the conformal primary basis, the soft photon theorems in (5) can be expressed as the following celestial correlation functions:
$\left\langle\mathcal{S}^{1}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{e_{a}}{w-z_{a}}\quad\left\langle\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,$ (15)
$\left\langle\mathcal{S}^{1-2j}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle={\frac{1}{(2j)!}}\sum_{a=1}^{N}\frac{e_{a}^{2j+1}\eta_{a}^{2j}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]$
$\times\Bigg[\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}e_{c}\,\frac{(w-z_{c})(\bar{w}-\bar{z}_{c})}{(z_{a}-z_{c})(\bar{z}_{a}-\bar{z}_{c})}\Bigg]^{2j-1}\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2j}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,$ (16)
for $j=1/2,1,3/2,\cdots$. In the above equations we have kept the helicity and all other internal quantum numbers of hard particles implicit because the soft factors depend only on the electric charges and the momenta of them. The soft operator $\mathcal{S}^{\Delta}(w,\bar{w})$ for positive helicity outgoing soft photon has been defined via the Mellin transformation of the photon annihilation operator $a(k,\sigma=+1)$ along with the conformal soft limit Donnay:2018neh taken inside the celestial correlation functions as [^7] Here we used the following result of Mellin transformation $\int_{0}^{\infty}d\omega\,\omega^{s-1}\omega^{n}(\ln\omega)^{k}=\frac{(-1)^{k}k!}{(s+n)^{k+1}}$, as derived in Krishna:2023ukw . One may be tempted to resum all the multi-order poles in the scaling dimensions and interpret the resummed Mellin-space Ward identity as a simple pole in $\Delta$ with anomalous dimensions. However, the resulting structure of the sum appears to be of the form $\sum_{n=0}^{\infty}(\Delta+n)^{n+2}$ in Mellin space, or equivalently exhibits the integral behavior $\int_{0}^{\infty}d\omega\omega^{\omega+\Delta-2}$, which falls into the category of the “Sophomore’s Dream”, for which no closed analytical result is known.
$\displaystyle\begin{gathered}\mathcal{S}^{1-2j}(w,\bar{w})=-\sqrt{2}(4\pi^{2})^{2j}\frac{(-1)^{2j}}{(2j)!}\lim_{\Delta\rightarrow 1-2j}(\Delta+2j-1)^{2j+1}\Gamma_{\Delta}^{+}(w,\bar{w}),\\ \Gamma_{\Delta}^{+}(w,\bar{w})=\int_{0}^{\infty}d\omega\,\omega^{\Delta-1}\ a(-\omega\,\mathbf{n}(w,\bar{w}),\sigma=+1),\end{gathered}$ (19)
for $j=0,1/2,1,3/2,\cdots$. The holomorphic and anti-holomorphic conformal dimensions of $\mathcal{S}^{1-2j}$ are $h=1-j,\,\bar{h}=-j$ as follows from the (Lorentz) conformal invariance of (15) and (16).
The Weinberg soft theorem (15) is known to be equivalent to the Ward-identity of a $U(1)$ Kac-Moody algebra at level zero. In the rest of the paper we focus on the symmetry interpretation of the all loop soft theorems (16).
IV $\mathcal{O}(\ln\omega)$ soft theorem as Ward identity and the dipole current operator
In the conformal primary basis, the $\mathcal{O}(\ln\omega)$ soft theorem (16) $(j=1/2)$ can be written as
$\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle$ (20)
This equation, as it stands, does not allow us to define a Celestial OPE between soft and hard operators because of the multi-particle sum. For example, if we take $w\rightarrow z_{a}$ then the most singular term in $(w-z_{a})$
$\mathcal{S}^{0}(w,\bar{w})\phi^{(\eta_{a})}_{\Delta_{a}}(z_{a},\bar{z}_{a},e_{a})\sim\frac{\eta_{a}e_{a}^{2}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\right)\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})$ (21)
does not have the standard form of an OPE. To cure this problem we introduce the “dipole current operator” $\bar{d}(\bar{w},\bar{z})$ whose correlation function is given by
$\boxed{\left\langle\bar{d}(\bar{w},\bar{z})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a}e_{a}\frac{\bar{w}-\bar{z}_{a}}{\bar{z}-\bar{z}_{a}}\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle}$ (22)
From (22) it is evident that the charge $Q_{\bar{d}}(\bar{w})$ defined through $Q_{\bar{d}}(\bar{w})=\oint\frac{d\bar{z}}{2\pi i}\bar{d}(\bar{w},\bar{z})$ measures the (anti-holomorphic) dipole moment of a charged operator, i.e,
$\displaystyle\left[Q_{\bar{d}}(\bar{w}),\phi^{(\eta)}_{\Delta}(z,\bar{z},e)\right]$ $=-e(\bar{z}-\bar{w})\phi^{(\eta)}_{\Delta}(z,\bar{z},e),$ (23)
justifying the name “dipole current operator”.
Now the $\mathcal{O}(\ln\omega)$ soft theorem can be rewritten as a Ward identity
$\boxed{\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}e_{a}^{2}}{w-z_{a}}\left\langle:\bar{d}\phi^{(\eta_{a})}_{\Delta_{a}-1}:(\bar{w},z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,}$ (24)
where the composite operator $\bar{d}\phi^{(\eta)}_{\Delta}:(\bar{w},z,\bar{z},e)$ is defined in the standard way as
$\displaystyle\begin{gathered}:\bar{d}\phi^{(\eta)}_{\Delta}:(\bar{w},z,\bar{z},e)\equiv\lim\limits_{\bar{\xi}\rightarrow\bar{z}}\Bigg(\bar{d}(\bar{w},\bar{\xi})\phi^{(\eta)}_{\Delta}(z,\bar{z},e)-e\frac{\bar{w}-\bar{z}}{\bar{\xi}-\bar{z}}\phi^{(\eta)}_{\Delta}(z,\bar{z},e)\Bigg).\end{gathered}$ (26)
The Ward identity (24) allows us to write down the Celestial OPE between the soft operator $\mathcal{S}^{0}$ and a hard operator as
$\boxed{\mathcal{S}^{0}(w,\bar{w})\phi^{(\eta)}_{\Delta}(z,\bar{z},e)\sim\frac{\eta e^{2}}{w-z}:\bar{d}\phi^{(\eta)}_{\Delta-1}:(\bar{w},z,\bar{z},e)\,.}$ (27)
The equation (22) implies that:
The dual Celestial CFT2 has a “Dipole Symmetry” generated by the current $\bar{d}(\bar{w},\bar{z})$. The existence of the dipole current follows from the locality of the celestial CFT2 in the presence of loop corrections in the bulk.
V Conformal transformation law of the dipole operator
Under the $SL(2,\mathbb{R})_{\text{R}}$ transformation $\bar{z}\rightarrow\bar{z}^{\prime}=\frac{a\bar{z}+b}{c\bar{z}+d}$ with $a,b,c,d\in\mathbb{R}$ and adbc=1ad-bc=1, the invariance of the correlation function (22) requires the dipole current $\bar{d}(\bar{w},\bar{z})$ to transform in the following way,
$\displaystyle\bar{d}(\bar{w},\bar{z})\rightarrow\bar{d}^{\prime}(\bar{w},\bar{z})=\frac{c\bar{w}+d}{c\bar{z}+d}\times\bar{d}\left(\bar{w}^{\prime},\bar{z}^{\prime}\right).$ (28)
Note that the RHS of the correlation function (22) is a polynomial of degree one in $\bar{w}$. So we can Taylor expand $\bar{d}(\bar{w},\bar{z})$ in $\bar{w}$ coordinate Banerjee:2020zlg ; Guevara:2021abz ; Strominger:2021mtt and write
$\bar{d}(\bar{w},\bar{z})=\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})+\bar{w}\ \bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}).$ (29)
In terms of the doublet $(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}},\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}})$ the transformation (28) becomes
$\displaystyle\bar{d}^{\prime\frac{1}{2}}_{\frac{1}{2}}(\bar{z})=\frac{1}{c\bar{z}+d}\left(a\,\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}^{\prime})+c\,\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z}^{\prime})\right),\quad\bar{d}^{\prime\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})=\frac{1}{c\bar{z}+d}\left(b\,\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}^{\prime})+d\,\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z}^{\prime})\right).$ (30)
and for infinitesimal translation, dilatation and special conformal transformations (SCT)888 $\displaystyle\text{Translation:}\quad\begin{pmatrix}1&b\\ 0&1\end{pmatrix},\qquad\text{Dilatation:}\quad\begin{pmatrix}\lambda^{\frac{1}{2}}&0\\ 0&\lambda^{-\frac{1}{2}}\end{pmatrix},\qquad\text{SCT:}\quad\begin{pmatrix}1&0\\ -c&1\end{pmatrix}$, (31) we get the following commutation relations
$\displaystyle\begin{gathered}\left[\bar{L}_{-1},\bar{d}^{j}_{m}(\bar{z})\right]=\bar{\partial}\bar{d}^{j}_{m}(\bar{z})+(j-m)\bar{d}^{j}_{m+1},\\ \left[\bar{L}_{0},\bar{d}^{j}_{m}(\bar{z})\right]=\left(j+m+\bar{z}\bar{\partial}\right)\bar{d}^{j}_{m}(\bar{z}),\\ \left[\bar{L}_{1},\bar{d}^{j}_{m}(\bar{z})\right]=\left(2j\bar{z}+\bar{z}^{2}\bar{\partial}\right)\bar{d}^{j}_{m}(\bar{z})-(j+m)\bar{d}^{j}_{m-1}(\bar{z}).\end{gathered}$ (35)
Here $j=\frac{1}{2}$, $m=\pm\frac{1}{2}$ and the generators of translation, dilatation and SCT are denoted by $\bar{L}_{-1},\,\bar{L}_{0}$ and $\bar{L}_{1}$ respectively.
The transformation law (30) shows that the doublet $\left(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z}),\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})\right)$ transform in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation with weight $\bar{h}=1/2$ and the other one is a two dimensional spin-$\frac{1}{2}$ representation. Under the spin-$\frac{1}{2}$ representation the fields $\bar{d}^{\frac{1}{2}}_{\pm\frac{1}{2}}(\bar{z})$ have weights $\pm\frac{1}{2}$. To see this at the level of infinitesimal transformation law, we denote the generator in the highest weight ($\bar{h}=\frac{1}{2}$) representation by $\bar{L}_{n}^{\bar{h}}$ and the generator in the spin-$\frac{1}{2}$ representation by $\bar{L}_{n}^{\bar{\sigma}}$. Now one can easily verify that the following commutation relations
$\displaystyle\begin{gathered}\left[\bar{L}^{\bar{h}}_{n},\bar{d}^{j}_{m}(\bar{z})\right]=\left(\bar{z}^{n+1}\bar{\partial}+(n+1)j\bar{z}^{n}\right)\bar{d}^{j}_{m}(\bar{z}),\\ \left[\bar{L}^{\bar{\sigma}}_{-1},\bar{d}^{j}_{m}(\bar{z})\right]=(j-m)\,\bar{d}^{j}_{m+1}(\bar{z}),\\ \left[\bar{L}^{\bar{\sigma}}_{0},\bar{d}^{j}_{m}(\bar{z})\right]=m\,\bar{d}^{j}_{m}(\bar{z}),\\ \left[\bar{L}^{\bar{\sigma}}_{1},\bar{d}^{j}_{m}(\bar{z})\right]=-(j+m)\,\bar{d}^{j}_{m-1}(\bar{z}),\end{gathered}$ (40)
reproduce the transformation law (35) for n=1,0,1n=-1,0,1, $j=\frac{1}{2}$ and $m=\pm\frac{1}{2}$ provided the relation $\bar{L}_{n}=\bar{L}_{n}^{\bar{h}}+\bar{L}_{n}^{\bar{\sigma}}$ holds.
VI Some notation
To make the $SL(2,\mathbb{R})_{\text{R}}$ covariance manifest it will be useful to introduce the two-component spinor notation. We define
$\bar{Z}^{\alpha}=\begin{pmatrix}\bar{z}\\ 1\end{pmatrix},\ \epsilon_{\alpha\beta}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},\ \bar{Z}_{\alpha}=\epsilon_{\alpha\beta}\bar{Z}^{\beta}=\begin{pmatrix}1\\ -\bar{z}\end{pmatrix},\ \alpha,\beta=\pm 1/2,$ (41)
so that (29) can be rewritten as
$\bar{d}(\bar{w},\bar{z})=\bar{W}^{\alpha}\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z}),$ (42)
and the Ward identity (22) for the dipole symmetry becomes [^9] Note that $\bar{z}-\bar{z}_{a}=\bar{Z}^{\alpha}\bar{Z}_{a\alpha}$, how ever we are keeping it explicit.
$\left\langle\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a}e_{a}\frac{\bar{Z}_{a\alpha}}{\bar{z}-\bar{z}_{a}}\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.$ (43)
We can also expand the soft photon operator $\mathcal{S}^{0}(w,\bar{w})$ in a similar way
$\mathcal{S}^{0}(w,\bar{w})=\bar{W}^{\alpha}\mathcal{S}^{0}_{\alpha}(w),$ (44)
and rewrite the $\mathcal{O}(\ln\omega)$ soft theorem as
$\displaystyle\begin{gathered}\left\langle\mathcal{S}^{0}_{\alpha}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\left\langle:\bar{d}^{\frac{1}{2}}_{\alpha}\phi^{(\eta_{a})}_{\Delta_{a}-1}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (46)
VII Correlation functions of dipole currents and matter fields
The dipole operator $\bar{d}^{\frac{1}{2}}_{\alpha}$ has scaling dimension $1/2$. Therefore conformal invariance determines the two point function to be  [^10] In terms of the original field $\bar{d}(\bar{w},\bar{z})$ this is nothing but $\langle\bar{d}(\bar{w}_{1},\bar{z}_{1})\bar{d}(\bar{w}_{2},\bar{z}_{2})\rangle=k\frac{\bar{w}_{1}-\bar{w}_{2}}{\bar{z}_{1}-\bar{z}_{2}}$. This form of the two point function is uniquely determined by the $SL(2,\mathbb{R})_{\text{R}}$ transformation law (28).
$\langle\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z}_{1})\bar{d}^{\frac{1}{2}}_{\beta}(\bar{z}_{2})\rangle=k\frac{\epsilon_{\alpha\beta}}{\bar{z}_{1}-\bar{z}_{2}}\,$, (47)
where $k$ is a real number which can be zero also. However, as we will see we can take $k$ to be nonzero without changing the single soft photon theorems. Our current understanding of the role of dipole symmetry in the loop level soft theorems does not allow us to determine the value of $k$. This is discussed in more detail in section XI.
The singular terms in the OPE between two dipole operators is given by
$\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z}_{1})\bar{d}^{\frac{1}{2}}_{\beta}({\bar{z}_{2}})\sim k\frac{\epsilon_{\alpha\beta}}{\bar{z}_{1}-\bar{z}_{2}}\,.$ (48)
This together with the boundary condition that in the absence of any operator insertion at infinity $\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z})\rightarrow 1/\bar{z}$ as $\bar{z}\rightarrow\infty$ determine the correlation function with multiple insertion of dipole currents to be
$\displaystyle\begin{gathered}\left\langle\bar{d}^{\frac{1}{2}}_{\alpha_{1}}(\bar{\xi}_{1})\prod_{p=2}^{n}\bar{d}^{\frac{1}{2}}_{\alpha_{p}}(\bar{\xi_{p}})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\left(\sum_{a}e_{a}\frac{\bar{Z}_{a{\alpha_{1}}}}{\bar{\xi_{1}}-\bar{z}_{a}}\right)\left\langle\prod_{p=2}^{n}\bar{d}^{\frac{1}{2}}_{\alpha_{p}}(\bar{\xi_{p}})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle\\ +\sum_{a=2}^{n}k\frac{\epsilon_{\alpha_{1}\alpha_{a}}}{\bar{\xi}_{1}-\bar{\xi}_{a}}\left\langle\bar{d}^{\frac{1}{2}}_{\alpha_{2}}(\xi_{2})\cdots\cancel{\bar{d}^{\frac{1}{2}}_{\alpha_{a}}(\bar{\xi}_{a})}\cdots\bar{d}^{\frac{1}{2}}_{\alpha_{n}}(\bar{\xi}_{n})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (51)
This correlation function allows us to define currents of different weights and spins and play an important role in this paper. For example, let us consider the correlation function
$\displaystyle\begin{gathered}\left\langle\bar{d}^{\frac{1}{2}}_{\alpha_{1}}(\bar{\xi}_{1})\bar{d}^{\frac{1}{2}}_{\alpha_{2}}(\bar{\xi}_{2})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\\ \left(\sum_{a}e_{a}\frac{\bar{Z}_{a{\alpha_{1}}}}{\bar{\xi_{1}}-\bar{z}_{a}}\right)\left(\sum_{b}e_{b}\frac{\bar{Z}_{b{\alpha_{2}}}}{\bar{\xi_{2}}-\bar{z}_{b}}\right)\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle+k\frac{\epsilon_{\alpha_{1}\alpha_{2}}}{\bar{\xi}_{1}-\bar{\xi}_{2}}\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,\end{gathered}$ (54)
with two insertions of the dipole current. This allows us to define a dimension $\bar{h}=1$ and spin-1 current $\bar{d}^{1}_{\alpha\beta}$ as
$\bar{d}^{1}_{\alpha\beta}(\bar{z})=\bar{d}^{1}_{\beta\alpha}(\bar{z})=\lim_{\bar{z}^{\prime}\rightarrow\bar{z}}\left(\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z}^{\prime})\bar{d}^{\frac{1}{2}}_{\beta}(\bar{z})-k\frac{\epsilon_{\alpha\beta}}{\bar{z}^{\prime}-\bar{z}}\right)=:\bar{d}^{\frac{1}{2}}_{\alpha}(\bar{z})\bar{d}^{\frac{1}{2}}_{\beta}(\bar{z}):\,$, (55)
whose correlation function with matter is given by
$\displaystyle\begin{gathered}\left\langle\bar{d}^{1}_{\alpha_{1}\alpha_{2}}(\bar{\xi})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\left(\sum_{a}e_{a}\frac{\bar{Z}_{a{\alpha_{1}}}}{\bar{\xi}-\bar{z}_{a}}\right)\left(\sum_{b}e_{b}\frac{\bar{Z}_{b{\alpha_{2}}}}{\bar{\xi}-\bar{z}_{b}}\right)\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (57)
In a similar fashion we can define spin-$j$ current $\bar{d}^{j}_{\alpha_{1}\cdots\alpha_{2j}}$ by generalizing the normal ordering (55) for $2j$ number of $\bar{d}^{\frac{1}{2}}$ in a standard way by removing all the singular terms and get
$\bar{d}^{j}_{\alpha_{1}\cdots\alpha_{2j}}=:\bar{d}^{\frac{1}{2}}_{\alpha_{1}}\bar{d}^{\frac{1}{2}}_{\alpha_{2}}\cdots\bar{d}^{\frac{1}{2}}_{\alpha_{2j}}:\,$, (58)
whose correlation function with matter is given by
$\displaystyle\begin{gathered}\left\langle\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{\xi})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\prod_{k=1}^{2j}\left(\sum_{a}e_{a}\frac{\bar{Z}_{a{\alpha_{k}}}}{\bar{\xi}-\bar{z}_{a}}\right)\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (60)
VIII $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ soft theorem as Ward identity
In this section, we focus on the singular parts of the soft theorems in (16) which contain poles in $w$. The Weinberg (15) and the singular parts of the logarithmic (16) soft theorems can be combined as
$\displaystyle\begin{gathered}\left\langle\mathcal{S}^{1-2j}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\frac{1}{(2j)!}\sum_{a=1}^{N}\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]^{2j}\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2j}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,\end{gathered}$ (63)
where $j=0,\frac{1}{2},1,\cdots$. Note that for $j=\frac{1}{2}$ the singular part coincides with the actual soft theorem whose Ward identity interpretation has been discussed in section IV. In this section we start with the analysis of the (singular part of the) soft theorem for $j=1$ using the notation introduced in section VI.
We write the $j=1$ soft photon $\mathcal{S}^{-1}(w,\bar{w})$ as
$\mathcal{S}^{-1}(w,\bar{w})=\frac{1}{2}\bar{W}^{\alpha}\bar{W}^{\beta}\mathcal{S}^{-1}_{\alpha\beta}(w).$ (64)
and rewrite the singular part of the $j=1$ soft theorem as
$\begin{gathered}\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\sum_{a=1}^{N}\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{Z}_{b\alpha}}{\bar{z}_{a}-\bar{z}_{b}}\right)\left(\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}e_{c}\,\frac{\bar{Z}_{c\beta}}{\bar{z}_{a}-\bar{z}_{c}}\right)\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (65)
Now to interpret (65) as (singular part of) a Ward identity we, by comparing with (57), introduce a spin-1 current $\bar{d}^{1}_{\alpha\beta}(\bar{z})$ such that
$\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}=\sum_{a=1}^{N}\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left\langle:\bar{d}^{1}_{\alpha\beta}\phi^{(\eta_{a})}_{\Delta_{a}-2}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,$ (66)
with the normal ordering definition
$\displaystyle\begin{gathered}:\bar{d}^{1}_{\alpha\beta}\phi^{(\eta)}_{\Delta}(z,\bar{z},e):\quad=\quad\lim_{\bar{z}^{\prime}\rightarrow\bar{z}}\Big[\bar{d}^{1}_{\alpha\beta}(\bar{z}^{\prime})\phi^{(\eta)}_{\Delta}(z,\bar{z},e)-e^{2}\frac{\bar{Z}_{\alpha}\bar{Z}_{\beta}}{(\bar{z}^{\prime}-\bar{z})^{2}}\phi^{(\eta)}_{\Delta}(z,\bar{z},e)\\ -e\frac{\bar{Z}_{\alpha}}{\bar{z}^{\prime}-\bar{z}}:\bar{d}^{\frac{1}{2}}_{\beta}\phi^{(\eta)}_{\Delta}:(z,\bar{z},e)-e\frac{\bar{Z}_{\beta}}{\bar{z}^{\prime}-\bar{z}}:\bar{d}^{\frac{1}{2}}_{\alpha}\phi^{(\eta)}_{\Delta}:(z,\bar{z},e)\Big].\end{gathered}$ (69)
Now generalization of the above construction to arbitrary values of $j$ is simple. We write the soft operator in spinor notation
$\mathcal{S}^{1-2j}(w,\bar{w})=\frac{1}{(2j)!}\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w),$ (70)
and using this we can rewrite the singular part (63) of the soft theorem as
$\begin{gathered}\left\langle\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\sum_{a=1}^{N}\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\left\langle:\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}\phi^{(\eta_{a})}_{\Delta_{a}-2j}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,\end{gathered}$ (71)
where the definition of the higher spin current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{z})$ is given in (58). The current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}$ is completely symmetric in the spinor indices and transforms in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation of weight $\bar{h}=j$ and the other one is the $(2j+1)$-dimensional spin-$j$ representation. We discuss this in more detail in section IX.1.
In the following section we discuss the algebra of these currents for general values of $k$.
IX Algebra of higher spin currents
Since the contraction between two dipole currents $\bar{d}^{\frac{1}{2}}$ is a $c$-number we can compute the OPE between the currents defined in (58) using Wick’s theorem. We show some results for the singular terms only:
$\bar{d}^{1/2}_{\alpha_{1}}(\bar{z}_{1})\bar{d}^{1/2}_{\alpha_{2}}(\bar{z}_{2})\sim k\frac{\epsilon_{\alpha_{1}\alpha_{2}}}{\bar{z}_{1}-\bar{z}_{2}}\,$, (72)
$\bar{d}^{1}_{\alpha_{1}\alpha_{2}}(\bar{z}_{1})\bar{d}^{1/2}_{\alpha_{3}}(\bar{z}_{2})\sim k\frac{\epsilon_{\alpha_{1}\alpha_{3}}\bar{d}^{1/2}_{\alpha_{2}}(\bar{z}_{2})+\epsilon_{\alpha_{2}\alpha_{3}}\bar{d}^{1/2}_{\alpha_{1}}(\bar{z}_{2})}{\bar{z}_{1}-\bar{z}_{2}}\,$, (73)
[…]
We can see that the currents $\bar{d}^{1}_{\alpha\beta}$ generate a $SL(2,\mathbb{R})$ Kac-Moody algebra. The currents $\bar{d}^{1/2}_{\alpha}$ and $\bar{d}^{1}_{\alpha\beta}$ form a closed subalgebra. However, starting from spin-$\frac{3}{2}$ the algebra of currents do not close among themselves. This is evident from the existence of the blue coloured terms in the OPE (83). To understand the underlying algebra better let us take the semiclassical limit of $k(\sim i\hbar)\rightarrow 0$. In this limit the problematic terms which are all of $\mathcal{O}(k^{2})$ and higher are subleading. The leading terms of $\mathcal{O}(k)$ in the OPE are those which arise from single contractions. If we keep only the leading terms then we get the closed OPE
$\begin{gathered}\bar{d}^{j_{1}}_{\alpha_{1}\cdots\alpha_{2j_{1}}}(\bar{z}_{1})\bar{d}^{j_{2}}_{\beta_{1}\cdots\beta_{2j_{2}}}(\bar{z}_{2})\sim k\ \frac{\epsilon_{\alpha_{1}\beta_{1}}\bar{d}^{j_{1}+j_{2}-1}_{\alpha_{2}\cdots\alpha_{2j_{1}}\beta_{2}\cdots\beta_{2j_{2}}}(\bar{z}_{2})+\text{Contractions}}{\bar{z}_{1}-\bar{z}_{2}}+\mathcal{O}(k^{2}).\end{gathered}$ (84)
Now we show that the global part of the algebra generated by the higher spin currents in the semiclassical limit is the wedge subalgebra of the $w_{1+\infty}$ algebra. We denote the wedge subalgebra by $\wedge w_{1+\infty}$.
We define the global modes
$\bar{\mathbf{d}}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}=\oint\frac{d\bar{z}}{2\pi i}\ \bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{z}).$ (85)
The algebra of these global modes is given by
$\begin{gathered}\left[\bar{\mathbf{d}}^{j_{1}}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j_{1}}},\bar{\mathbf{d}}^{j_{2}}_{\beta_{1}\beta_{2}\cdots\beta_{2j_{2}}}\right]=k\ \left(\epsilon_{\alpha_{1}\beta_{1}}\bar{\mathbf{d}}^{j_{1}+j_{2}-1}_{\alpha_{2}\cdots\alpha_{2j_{1}}\beta_{2}\cdots\beta_{2j_{2}}}+\cdots+\epsilon_{\alpha_{2j_{1}}\beta_{2j_{2}}}\bar{\mathbf{d}}^{j_{1}+j_{2}-1}_{\alpha_{1}\cdots\alpha_{2j_{1}-1}\beta_{1}\cdots\beta_{2j_{2}-1}}\right),\end{gathered}$ (86)
containing $2j_{1}\times 2j_{2}$ number of terms in the RHS. To show that this algebra is isomorphic to $\wedge w_{1+\infty}$ let us define
$w^{p}_{n}\equiv\frac{1}{2}\bar{\mathbf{d}}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}},$ (87)
where
$p=j+1,\ n=\sum_{i=1}^{2j}\alpha_{i},\ \alpha_{i}=\pm\frac{1}{2},\ j\in\frac{1}{2}\mathbb{Z_{+}}\,.$ (88)
It follows from (88) that the maximum value of $n$ is j=p1j=p-1 and the minimum value of $n$ is j=1p-j=1-p. In other words
$1-p\leq n\leq p-1.$ (89)
It is easy to show that in terms of the generators $w^{p}_{n}$ the algebra (86) can be rewritten as
$\left[w^{p}_{m},w^{q}_{n}\right]=k\left[m(q-1)-n(p-1)\right]w^{p+q-2}_{m+n},$ (90)
with $p,q=\frac{3}{2},2,\frac{5}{2},3,\cdots$ and the value for $w^{1}_{0}=\frac{1}{2}$ appears when $p=q=\frac{3}{2}$.
The algebra (90) subject to the restriction (89) is well known to be the wedge subalgebra $\wedge w_{1+\infty}$ Bakas:1989xu ; Pope:1991zka ; Pope:1991ig . This is a classical algebra because we have neglected all the terms involving more than one contraction. If $k=0$ then of course this reduces to an infinite dimensional Abelian algebra. In the context of tree level soft graviton theorems $w_{1+\infty}$ has already appeared as a symmetry of the dual Celestial CFT2 Strominger:2021mtt ; Adamo:2021lrv .
IX.1 Conformal transformation of higher spin currents
We start with the current $\bar{d}^{j}$ defined in (58), which appears in the singular part of the Ward identity for the $\mathcal{O}\left(\omega^{2j-1}(\ln\omega)^{2j}\right)$ soft theorem in (71). Contracting this current with $2j$ insertions of $\bar{W}$, we obtain
$\displaystyle\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\,\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{z})$ $=$ $\sum_{m=-j}^{j}{}^{2j}C_{j+m}\bar{w}^{j+m}\,:\left[\left(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z})\right)^{j+m}\left(\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})\right)^{j-m}\right]:$ (91)
$\equiv$ $\sum_{m=-j}^{j}{}^{2j}C_{j+m}\quad\bar{d}^{j}_{m}(\bar{z})\,\bar{w}^{j+m},$
where the last line defines $\bar{d}^{j}_{m}(\bar{z})$ with $m=\sum\limits_{i=1}^{2j}\alpha_{i}$ in relation to the LHS, and it’s explicit expression is given by
$\bar{d}^{j}_{m}(\bar{z})\equiv\quad:\left[\left(\bar{d}^{\frac{1}{2}}_{\frac{1}{2}}(\bar{z})\right)^{j+m}\left(\bar{d}^{\frac{1}{2}}_{-\frac{1}{2}}(\bar{z})\right)^{j-m}\right]:\,$, (92)
for $j\leq m\leq j$. In relation to the dipole operator $\bar{d}(\bar{w},\bar{z})$ with correlation function (22), the RHS of (91) can also be identified as normal ordering of $2j$ number of dipole operator, i.e.
$\sum_{m=-j}^{j}{}^{2j}C_{j+m}\quad\bar{d}^{j}_{m}(\bar{z})\,\bar{w}^{j+m}=:\left[\bar{d}(\bar{w},\bar{z})\right]^{2j}:\,.$ (93)
Now under the $SL(2,\mathbb{R})_{\text{R}}$ transformation described in section V, we find the following transformation rule
$\left[\bar{d}^{\prime}(\bar{w},\bar{z})\right]^{2j}:=\frac{(c\bar{w}+d)^{2j}}{(c\bar{z}+d)^{2j}}:\left[\bar{d}(\bar{w}^{\prime},\bar{z}^{\prime})\right]^{2j}:\,.$ (94)
This in turn provides the following relation for transformation of $\bar{d}^{j}_{m}$:
$\displaystyle\sum_{m=-j}^{j}{}^{2j}C_{j+m}\bar{w}^{j+m}\bar{d}^{\prime j}_{m}(\bar{z})=\frac{(c\bar{w}+d)^{2j}}{(c\bar{z}+d)^{2j}}\sum_{n=-j}^{j}{}^{2j}C_{j+n}\left(\frac{a\bar{w}+b}{c\bar{w}+d}\right)^{j+n}\bar{d}^{j}_{n}\left(\frac{a\bar{z}+b}{c\bar{z}+d}\right).$ (95)
Now focusing on infinitesimal $SL(2,\mathbb{R})_{\text{R}}$ transformations as in section V we find the following commutation relations
$\displaystyle\begin{gathered}\left[\bar{L}_{-1},\bar{d}^{j}_{m}(\bar{z})\right]=\bar{\partial}\bar{d}^{j}_{m}(\bar{z})+(j-m)\bar{d}^{j}_{m+1},\\ \left[\bar{L}_{0},\bar{d}^{j}_{m}(\bar{z})\right]=\left(j+m+\bar{z}\bar{\partial}\right)\bar{d}^{j}_{m}(\bar{z}),\\ \left[\bar{L}_{1},\bar{d}^{j}_{m}(\bar{z})\right]=\left(2j\bar{z}+\bar{z}^{2}\bar{\partial}\right)\bar{d}^{j}_{m}(\bar{z})-(j+m)\bar{d}^{j}_{m-1}(\bar{z}),\end{gathered}$ (99)
for $j\in\frac{1}{2}\mathbb{Z}_{+}$ and $j\leq m\leq j$. This coincides with the algebra given in (35) for $j=\frac{1}{2}$.
X Modifications of currents at finite $k$ and their impact on single soft theorems
As we pointed out earlier, the current algebra involving $\bar{d}^{j}$ for $j\geq\frac{3}{2}$ does not close among themselves at the quantum level. For example, this is evident from the presence of the blue terms in (83) for the OPE of two spin-$\frac{3}{2}$ currents. Thus, at the quantum level, in order to demand a closed algebra among currents of spin $j\geq\frac{3}{2}$, we need to modify the definitions of currents in (58) in the following way:
$\bar{D}^{\frac{3}{2}}{\alpha_{1}\alpha_{2}\alpha_{3}}=\bar{d}^{\frac{3}{2}}_{\alpha_{1}\alpha_{2}\alpha_{3}}+kA\left(\epsilon_{\alpha_{1}\alpha_{2}}\bar{\partial}\bar{d}^{\frac{1}{2}}_{\alpha_{3}}+\text{symmetrization}\right),$ (100)
$\begin{gathered}\bar{D}^{2}{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}=\bar{d}^{2}_{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}+kB\left(\epsilon_{\alpha_{1}\alpha_{2}}:\bar{d}^{\frac{1}{2}}_{\alpha_{3}}\bar{\partial}\bar{d}^{\frac{1}{2}}_{\alpha_{4}}:+\text{symmetrization}\right)\\ +kC\left(\epsilon_{\alpha_{1}\alpha_{2}}\bar{\partial}\bar{d}^{1}_{\alpha_{3}\alpha_{4}}+\text{symmetrization}\right)\end{gathered}$ (101)
where $A,B,C$ are dimensionless numbers independent of $k$ and are purely quantum corrections since $k\sim i\hbar$. Possibly they can be fixed by demanding that the algebra among $\bar{D}^{j}$ closes along the same line as in Bergshoeff:1991un . Now the question is whether, for non-vanishing $A,B,C$ the soft theorems receive modifications starting at the 3-loop order if we replace $\bar{d}^{j}$ by $\bar{D}^{j}$ in (71). The answer is no: the soft theorem Ward identities in (16) do not receive corrections as the $\bar{D}^{j}$s appear contracted with $2j$ numbers of $\bar{W}$ and $\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\bar{D}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}=\bar{W}^{\alpha_{1}}\bar{W}^{\alpha_{2}}\cdots\bar{W}^{\alpha_{2j}}\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}$ even when $A,B,C\cdots$ above are non-vanishing. So there is no way to read off the values of $A,B,C\cdots$ from single soft photon theorem expressions.
XI What is the value of $k$?
The reader must have realized the importance of the value of $k$. If $k$ is zero then the algebra of the higher spin current becomes Abelian. However, if $k$ is nonzero but small such that we can neglect the quantum corrections then the algebra of the higher spin currents becomes the well known $\wedge w_{1+\infty}$. Now in order to know the value of $k$ we need to calculate celestial correlators or $S$-matrix elements of the form
$\left\langle\bar{d}^{j_{1}}(\bar{w}_{1})\bar{d}^{j_{2}}(\bar{w}_{2})\prod_{i=1}^{N}\phi_{\Delta_{i}}^{(\eta_{i})}(z_{i},\bar{z}_{i},e_{i})\right\rangle.$ (102)
However, we do not know the Feynman diagram representation of the $S$-matrix elements of the form (102) because the higher spin currents do not appear as asymptotic states in a scattering experiment in an obvious way.
Now one may think that the (consecutive) double soft theorem is a way out. But, the problem is that the currents appear in the soft theorems only as composite operators of the form $\bar{d}^{j}(\bar{z})\phi_{\Delta}^{(\eta)}(z,\bar{z},e):$. If we want to derive the consecutive double soft theorem using the Celestial CFT then we encounter correlation functions with multiple insertions of the composite operators, for example,
$\left\langle:\bar{d}^{j_{1}}(\bar{z}_{1})\phi_{\Delta_{1}}^{(\eta_{1})}(z_{1},\bar{z}_{1},e_{1}):\ :\bar{d}^{j_{2}}(\bar{z}_{2})\phi_{\Delta_{2}}^{(\eta_{2})}(z_{2},\bar{z}_{2},e_{2}):\ \prod_{i=3}^{N}\phi_{\Delta_{i}}^{(\eta_{i})}(z_{i},\bar{z}_{i},e_{i})\right\rangle.$ (103)
This correlation function depends on $k$ but to compute it we also need the celestial OPE between the matter primaries $\phi_{\Delta_{1}}^{(\eta_{1})}(z_{1},\bar{z}_{1},e_{1})$ and $\phi_{\Delta_{2}}^{(\eta_{2})}(z_{2},\bar{z}_{2},e_{2})$. This of course depends on the theory under consideration. Therefore the value of $k$ may differ from theory to theory or it may be zero for every theory. We leave a more decisive statement on the value of $k$ to future research.
Acknowledgment
We would like to thank Alok Laddha, Ana-Maria Raclariu and Ashoke Sen for very helpful discussions. SB would also like to thank all the participants of the Pre-ISM Workshop in TIFR Mumbai, where this work was presented, for interesting questions and comments. We acknowledge the use of ChatGPT for providing general information and for assisting with exploratory checks to arrive at the projection operator (6). The work of SB is supported by the Swarnajayanti Fellowship (File No- SB/SJF/2021-22/14) of the Department of Science and Technology and ANRF, India. The work of RM is supported by ANRF-CRG project (CRG/2022/006165). The work of BS was supported by STFC under grant ST/X000753/1.
Appendix A Support for the correctness of conjectured $n$-loop soft factor
Based on the soft photon theorem derivation at order $\omega(\ln\omega)^{2}$ in Sahoo:2020ryf from the analysis of two-loop Feynman diagrams in four dimensional scalar QED in presence of scalar contact interaction, we find that for scattering of massless charged particles the only diagrams giving non-vanishing contributions are those in which one end of both the virtual photon propagators are attached to the charged particle line from which the soft photon is emitted. Rest of the diagrams give vanishing contributions to $\mathcal{O}\left(\omega(\ln\omega)^{2}\right)$ soft factor, when charged particles are massless. [^11] Note that massless charged particles can themselves mediate long-range interactions and may generate additional contributions to logarithmic soft photon factors. This can be readily verified by replacing one or more virtual photon lines in Fig. 1 with massless charged scalar propagators, allowing for the corresponding three-point interaction vertices among them. However, if interaction vertices involving three massless charged particles are disallowed, one can argue using power counting that the coefficients $\mathbb{S}^{(n)}_{\text{em}}$ are not modified in the absence of any other massless particles besides the photon and charged particles. After examining a few higher-loop orders, this feature appears to be generic for massless charged particle scattering. Now assuming this observation to be valid for all-loop order S-matrix in massless scalar QED, we find that the only $n$-loop Feynman diagrams in Fig.1 provide non-vanishing contributions to $\mathbb{S}_{\text{em}}^{(n-1)}$, when evaluated with G-photon virtual propagators Sahoo:2018lxl ; Sahoo:2020ryf . The first diagram in Fig.1 represents the set of diagrams in which all the $n$ virtual photon (blue wavy lines) propagators are attached between particle-$a$ and $n$ different charged particles (black solid lines) through scalar–scalar–photon vertices, and the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon vertex. The second diagram in Fig.1 represents the set of diagrams in which all but one of the $n$ virtual photon propagators are attached between particle-$a$ and $n$ different charged particles through scalar–scalar–photon vertices, while the soft photon is attached to particle-$a$ at the far right through a scalar–scalar–photon–photon vertex where one of the remaining virtual photon is also attached. The third diagram in Fig.1 represents the set of diagrams that differ from the first diagram only in that two or more virtual photons are now attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations. Similarly, the fourth diagram in Fig.1 represents the set of diagrams that differ from the second diagram only in that two or more virtual photons are attached between particle-$a$ and particle-$b_{i}$ via ladder-type or cross-ladder-type configurations.
$p_{b_{n+1}}$$p_{b_{N-1}}$$p_{b_{1}}$$p_{b_{2}}$$p_{a}$$p_{b_{3}}$$p_{b_{n}}$}$k$
$p_{b_{n+1}}$$p_{b_{N-1}}$$p_{b_{1}}$$p_{b_{2}}$$p_{a}$$p_{b_{3}}$$p_{b_{n}}$}$k$
$p_{b_{n-1}}$$p_{b_{N-1}}$$p_{b_{1}}$$p_{b_{2}}$$p_{a}$$p_{b_{3}}$$p_{b_{n-2}}$}$k$
$p_{b_{n-1}}$$p_{b_{N-1}}$$p_{b_{1}}$$p_{b_{2}}$$p_{a}$$p_{b_{3}}$$p_{b_{n-2}}$}$k$
Figure 1: $n$-loop Feynman diagrams contributing to $\mathbb{S}_{\text{em}}^{(n-1)}$ in the theory of massless scalar QED.
From the ratio of two amplitudes as follows from (1), the leading-log $n$-loop soft factor from the set of the Feynman diagrams in Fig.1 turns out to be
$\displaystyle S^{n-\text{loop}}_{\text{em}}\left(\{e_{a},p_{a}\},(\varepsilon^{h},k)\right)$ $=$ $\frac{1}{n!}\sum_{a=1}^{N}e_{a}\frac{\varepsilon^{h}_{\mu}k_{\nu}}{p_{a}.k}\left(p_{a}^{\mu}\frac{\partial K_{\text{em}}}{\partial p_{a\nu}}-p_{a}^{\nu}\frac{\partial K_{\text{em}}}{\partial p_{a\mu}}\right)\left(k_{\sigma}\frac{\partial K_{\text{em}}}{\partial p_{a\sigma}}\right)^{n-1},$ (104)
where the contributions of first term inside the parenthesis comes from the sum over first and third Feynman diagrams in Fig.1 and the contributions of second term inside the parenthesis comes from the sum over third and fourth Feynman diagrams in Fig.1. The expression for $K_{\text{em}}$ is given by
$\displaystyle K_{\text{em}}$ $=$ $\frac{i}{2}\sum_{\begin{subarray}{c}b,c=1\\ c\neq b\end{subarray}}^{N}e_{b}e_{c}\,(p_{b}.p_{c})\int_{\omega}^{\Lambda}\frac{d^{4}\ell}{(2\pi)^{4}}\frac{1}{(\ell^{2}-i\epsilon)(p_{b}.\ell+i\epsilon)(p_{c}.\ell-i\epsilon)}$ (105)
$\simeq$ $\frac{i}{2}(\ln\omega)\sum_{\begin{subarray}{c}b,c=1\\ c\neq b\end{subarray}}^{N}\frac{e_{b}e_{c}}{4\pi}\left[\delta_{\eta_{b}\eta_{c},1}+\frac{i}{2\pi}\log\left((p_{b}.p_{c})^{2}\right)\right],$
where $\Lambda$ is of order $\sqrt{|p_{b}.p_{c}|}$, and in the second line after using the sign $\simeq$ we only kept the $\mathcal{O}(\ln\omega)$ contribution. Note that, while evaluating the above integral following Sahoo:2018lxl , one encounters collinear divergences if the computation is performed directly in the massless case. However, all such collinearly divergent terms cancel once the sum over the particle indices $b,c$ is performed. Hence the massless limit of the integration result provided Sahoo:2018lxl for massive charged particles does not suffer from any collinear divergences. Following the terminology of Sahoo:2018lxl , we refer to the first term inside the parentheses, namely $\delta_{\eta_{b}\eta_{c},1}$, as the “classical” contribution, and the second term as the “quantum” contribution. For a careful analysis of the “classical” part of the above integral directly in the massless case, the reader is referred to the appendix of Choi:2025mzg . Now substituting the expression for $K_{\text{em}}$ in (104) we find the following soft factor coefficient at order $\omega^{n-1}(\ln\omega)^{n}$:
$\displaystyle\mathbb{S}_{\text{em}}^{(n-1)}=-\frac{1}{n!(4\pi^{2})^{n}}\sum_{\begin{subarray}{c}a,b=1\\ b\neq a\end{subarray}}^{N}\frac{e^{2}_{a}e_{b}}{p_{a}.p_{b}}\frac{\varepsilon^{h}_{\mu}\mathbf{n}_{\nu}}{p_{a}.\mathbf{n}}\left(p_{a}^{\mu}p_{b}^{\nu}-p_{a}^{\nu}p_{b}^{\mu}\right)\left(\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}\frac{e_{a}e_{c}}{p_{a}.p_{c}}(p_{c}.\mathbf{n})\right)^{n-1},$ (106)
which is quoted in (5).